OpenStudy (anonymous):

Please help me with this! Thank you! Find the amount owed at the end of 6 years if $2500 is loaned at a rate of 7% compounded quarterly.

OpenStudy (anonymous):

a. $3302.47 b. $3791.11 c. $4398.34 d. $3175.95

OpenStudy (amistre64):

id just use the compound formula:\[A=P(1+r/n)^{nt}\] P is the principal amount r is the rate n is the number of times compounded in a period t is the number of periods a period is usually 1 year so ......

OpenStudy (anonymous):

\[A=P \left( 1+\frac{ r }{100 } \right)^t\] t=6 years=4*6=24 quarters r=7per annum=7/4 per quarter. calculate A,given P=$2500

OpenStudy (anonymous):

Okay, thanks for the help!

OpenStudy (anonymous):

yw